Solve by factoring: a2 -5a-14 = 0
Answer by ewatrrr(24785)

 
Hi
a^2 -5a-14 = 0
(a - 7)(a + 2) = 0
(a - 7) = 0
(x + 2) = 0, a = -2, 7
F First terms
O Outside terms 2a
I Inside terms -7a
L Last terms